# Verifiable cloud-based variational quantum algorithms

### v2

@@ III The Proposed Protocol
We adopt the strategy outlined in Sueki et al. (2013), where the server performs measurements after coupling the ancillary qubits with the register qubits, effectively addressing the limitations of using Bell pairs to tolerate channel loss in Shingu et al.’s protocol. The client is required to perform measurements in the bases {|±$\left\{\ket{\pm}^{\phi_{k}}\mid\phi_{k}=\frac{k\pi}{4},\,k\in\{0,1,\ldots,7\}\right\}$. For verification, we use trap qubits to create dummy gates in the trap wires as the circuit is transformed into the universal gate patterns shown in Fig. 4. Additionally, we delegate encrypted measurements of the output register qubits to the server, making it difficult to identify the trap wires in the circuits.
The client delegates the $J(\phi)$ operators within the universal gate patterns shown in Fig. 5, while the server provides the $CZ$ gates. This process includes preparation, computation, verification, and parameter updating phases, as briefly illustrated in Fig. 6. The specific steps are as follows:
Figure 4: (a) The gate pattern circuit, which includes 8 $J(\phi)$ operators and 2 $CZ$ gates. (b) The universal gate patterns, which are composed of multiple blocks, each block representing a gate pattern.
@@
Figure 5: Realization of the $J(\phi)$ operator: The server sends one qubit from a Bell pair to the client, who then applies a $T$ gate $k$ times, where $k$ is randomly and uniformly selected from $\{0,1,\ldots,7\}$, followed by a Hadamard gate $H$, and measures in the $Z$ basis. This measurement is equivalent to measuring in the basis {|±$\left\{\ket{\pm}^{\frac{k\pi}{4}}\right\}$. The server then performs a sequence of operations on the ancillary and register qubits, including a fixed coupling operation $CX_{RA}(H_{R}\otimes I_{A})$, where $CX$ is the controlled-X gate and $I$ is the identity gate, followed by measurements on the ancillary qubits.
@@
Figure 6: Brief process of the proposed protocol: The server sends an ancillary qubit, as one half of a Bell pair, to the client, who measures it in the basis {|±$\left\{\ket{\pm}^{\frac{k\pi}{4}}\right\}$. The client then sends the reception status back to the server, requesting a resend if the qubit is lost. The server performs encrypted measurements as instructed by the client and returns the results, enabling the client to verify the server’s honesty and calculate encrypted measurement angles and gradients for the optimizer.
@@
  2. 2.
    The server prepares $N$ register qubits |$\ket{0}_{R}$, two ancillary qubits $\frac{1}{\sqrt{2}}(\ket{00}_{A}+\ket{11}_{A})$ as a Bell pair, one ancillary qubit |$\ket{+}_{A}$, and one ancillary qubit |$\ket{0}_{A}$. Meanwhile, the client chooses $\frac{2N}{3}$ register qubits as trap qubits for verification, which is optimal Morimae (2014).
@@
  1. 1.
    The server sends one qubit from the Bell pair to the client, who measures it in the basis {|±$\left\{\ket{\pm}^{\frac{k\pi}{4}}\right\}$, yielding the result $s_{0}$, where $k$ is randomly and uniformly selected from $\{0,1,\ldots,7\}$. The client generates a random bit $r$ to indicate the reception status, setting it to 0 if the particle arrived successfully or to 1 if it was lost. This status is communicated to the server, who generates a new Bell pair and resends half of it to the client until the qubit arrives successfully. The other half of the Bell pair will eventually be in the state $Z^{s_{0}}\ket{+}^{\frac{k\pi}{4}}_{A}$.
  2. 2.
    The server employs the transformed circuits $\{U^{(c)}_{\rm{AN}}\}_{c=1}^{G}$ using the $J(\phi)$ operator, as depicted in Fig. 5, to generate the trial wave functions $\left\{\left|\psi^{(c)}(\vec{\theta}[1])\right\rangle\right\}_{c=1}^{G}$. In each delegation of the $J(\phi)$ operator, the $Z$ measurement results $s_{1}$ and $s_{2}$ are first sent to the client. The client then uses $s_{1}$ to compute the encrypted measurement angle $\phi^{\prime}=\phi-(-1)^{s_{1}}\left(\frac{k\pi}{4}\right)+(r+s_{0})\pi$, where $r\in\{0,1\}$ is a random bit chosen by the client. This encrypted angle $\phi^{\prime}$ is sent to the server, which measures in the basis {|±$\left\{\ket{\pm}^{\phi^{\prime}}\right\}$. The resulting measurement outcome $s_{3}$ is then sent to the client. The actual measurement angle $\phi$ is adaptively chosen based on prior Pauli byproducts to ensure determinism in MBQC.
  3. 3.
    In each transformed circuit, a random selection of $\frac{2N}{3}$ wires is designated as trap wires, where dummy gates are implemented. On half of these trap wires, the identity gate is applied $M$ times on each wire. On the remaining trap wires, a single Hadamard gate is randomly inserted, with the identity gate occupying the other $M-1$ positions on each wire.
  4. 4.
    The server measures the output register qubits using the encrypted measurement angle $\phi_{\text{out}}^{{}^{\prime}}$ provided by the client, where $\phi_{\text{out}}^{{}^{\prime}}=(-1)^{x_{\text{out}}}\phi_{\text{out}}+z_{\text{out}}\pi$, with $\phi_{\text{out}}$ being the actual measurement angle. The measurement results are then sent to the client. The client sets $\phi_{out}$ to select the desired measurement basis in the $XY$plane. To measure in the $Z$ basis, the client delegates an additional $R_{y}\left(\frac{\pi}{2}\right)$ gate at the end of the circuit and sets $\phi_{out}=\pi$.
  5. 5.
@@
  6. 6.
    The client compensates for the Pauli byproduct effect by flipping the measurement result of the output register qubits.
@@
  1. 1.
    At the end of B2(3), the output register qubits consist of non-trap qubits and trap qubits in a random permutation designed by the client. The output state can be written as $\sigma_{q}P|\Psi\rangle_{R}=\sigma_{q}P\left(\ket{\psi}_{out}^{\frac{N}{3}}\otimes\ket{0}_{T}^{\frac{N}{3}}\otimes\ket{+}_{T}^{\frac{N}{3}}\right)$, where |$\ket{\psi}_{out}$ represents the non-trap qubits, and $\sigma_{q}$ is the Pauli byproduct operator. The subscript $R$ includes both non-trap qubits ($out$) and trap qubits ($T$), with $P$ as the permutation.
  2. 2.
    The client instructs the server to measure all desired trap qubits, |$\ket{0}_{T}^{\frac{N}{3}}$ in the $Z$ basis and |$\ket{+}_{T}^{\frac{N}{3}}$ in the $X$ basis, in B2(4). If any undesired output |$\ket{1}_{T}$ in the $Z$ basis or |$\ket{-}_{T}$ in the $X$ basis is obtained after compensating for the Pauli byproduct effect, the protocol is terminated.
@@
  1. 1.
    The server and the client repeat B2 to B3 $\{\mathcal{R}^{(c)}\}_{c=1}^{G}$ times to derive the expectation values of $\{\hat{A}^{(c)}_{1},\hat{A}^{(c)}_{2},\cdots,\hat{A}^{(c)}_{K^{(c)}}\}_{c=1}^{G}$.
  2. 2.
    The client computes the measurement angles for adopting circuits $\{U^{(c)}_{\rm{AN}}\}_{c=1}^{G}$ with $\vec{\theta}[2]$.
  3. 3.
    The client utilizes its optimizer to update the parameters, resulting in $\vec{\theta}[2]=(\theta_{1}[2],\cdots,\theta_{L}[2])^{T}$ for the next step.
  4. 4.
    The client and server reiterate the above steps $(\mathcal{I}-2)$ times with $\{U^{(c)}_{\rm{AN}}\}_{c=1}^{G}$ and $\vec{\theta}[j]$. Based on the results from the $j$-th step, classical computations update the client’s parameters to $\vec{\theta}[j+1]$ for $j=2,3,\ldots,\mathcal{I}-1$.
IV Analysis
IV.1 Verifiability
We use trap qubits to detect malicious operations on the output register qubits. The verifiability of the proposed protocol is demonstrated in Theorem I, which follows a method similar to that in Morimae (2014), where further details can be found.
Theorem 1: The probability of the client being tricked by the server is exponentially small.
@@ IV.1 Verifiability
If the server is malicious, it may deviate from the state $\rho=\sigma_{q}|\Psi\rangle_{R}\langle\Psi|_{R}\sigma^{\dagger}_{q}$ to any other state. However, due to the completely-positive-trace-preserving (CPTP) map Morimae and Fujii (2013b), this deviation can be detected as a random Pauli attack.
@@
We can calculate the probabilities that each operator in $\sigma_{\alpha}$ does not change any trap qubits. Suppose $\max(a,b,c)=a$. An $X$ operator that does not change any trap qubits will only act on |$\ket{+}_{T}$ and non-trap qubits, the number of which is $\frac{2N}{3}$. Thus, the probability is:
@@
The proposed protocol requires the client to send qubits, qubit reception statuses, and encrypted measurement angles, which violates the no-signaling principle. However, the permutation $P$ is not included in these transmissions, ensuring $P$ remains secret. By encrypting the measurement angles for the trap qubits after the circuit construction is completed, the server cannot distinguish between trap and non-trap qubits, further protecting $P$.
After the client randomly selects a permutation $P$, the probability that $P^{\dagger}\sigma_{\alpha}P$ does not alter any trap qubits is at most $\left(\frac{2}{3}\right)^{\frac{|\alpha|}{3}}$. Consequently, the probability that the client is deceived by the server is at most $\prod_{c=1}^{G}\left(\frac{2}{3}\right)^{\frac{\mathcal{R}^{(c)}|\alpha|}{3}}=\left(\frac{2}{3}\right)^{\frac{\sum_{c=1}^{G}\mathcal{R}^{(c)}|\alpha|}{3}}$, where $G$ is the number of original and variant quantum circuits used to calculate gradients, and $\{\mathcal{R}^{(c)}\}_{c=1}^{G}$ is the set of circuit repetitions. This probability becomes exponentially small when $\sum_{c=1}^{G}\mathcal{R}^{(c)}$ is sufficiently large, ensuring the protocol’s verifiability.
IV.2 Blindness and Correctness
@@
Theorem 2: The proposed protocol guarantees input, output, and algorithm blindness.
@@
Input Blindness: After the client measures one qubit of the Bell pair, the remaining qubit is left in a maximally mixed state, $\frac{1}{16}\sum_{s_{0}=0}^{1}\sum_{k=0}^{7}\left(Z^{s_{0}}\ket{+}^{\frac{k\pi}{4}}_{A}\bra{+}^{\frac{k\pi}{4}}_{A}Z^{s_{0}\dagger}\right)=\frac{I}{2}$, with the value of $k$ hidden from the server.
Output Blindness: After applying the $J(\phi)$ operator, the register qubit is $X^{s_{0}\oplus s_{3}\oplus r}HR_{z}(\phi)\ket{\psi}_{R}$. Upon completion of the delegated circuit, the output register qubits are $X^{x_{out}}Z^{z_{out}}U\ket{\Psi}_{R}$, where $X^{x_{out}}Z^{z_{out}}$ are Pauli byproducts. Since the server cannot determine $r$ and $s_{0}$, it cannot compensate for $X^{s_{0}\oplus s_{3}\oplus r}$, and therefore cannot compensate for $X^{x_{out}}Z^{z_{out}}$, leaving the output qubits in a maximally mixed state.
Algorithm Blindness: When the client sends the encrypted measurement angle $\phi^{\prime}=\phi-(-1)^{s_{1}}\frac{k\pi}{4}+(s_{0}+r)\pi$ to the server, the true measurement angle $\phi$ remains hidden from the server, as $k$ and $r$ are randomly chosen, and $s_{0}$ is never disclosed to the server.
@@
Theorem 3: The proposed protocol ensures correctness throughout the computation.
@@
In each gate pattern, as illustrated in Fig. 4(a), the server directly applies the $CZ$ gate, while the $J(\phi)$ operator is realized using the circuit shown in Fig. 5 from Sueki et al. (2013), resulting in $HR_{z}(\phi)\ket{\psi}_{R}$ with Pauli byproducts. The client adjusts the measurement results sent by the server based on these Pauli byproducts, ensuring the correct $J(\phi)$ operators in the universal gate patterns.
Arbitrary single-qubit gates can be implemented within this gate pattern. For two arbitrary single-qubit gates $U_{1}$ and $U_{2}$, we decompose $U_{1}$ as $R_{z}(\phi_{0})R_{x}(\phi_{1})R_{z}(\phi_{2})$ and $U_{2}$ as $R_{z}(\phi_{4})R_{x}(\phi_{5})R_{z}(\phi_{6})$ using Euler’s rotation theorem Gothen and Guedes de Oliveira (2023). This can be realized in the gate pattern as follows:
$\begin{split}U_{1}\otimes U_{2}&=\left(J_{1}(\phi_{0})J_{1}(\phi_{1})\right)\otimes\left(J_{2}(\phi_{4})J_{2}(\phi_{5}=0)\right)CZ_{12}\\ &\quad\left(J_{1}(\phi_{2})J_{1}(\phi_{3})\right)\otimes\left(J_{2}(\phi_{6})J_{2}(\phi_{7}=0)\right)CZ_{12}.\end{split}$ (4)
The $CX$ gate can also be implemented within this pattern:
$\begin{split}CX_{12}&=\left(J_{1}(\phi_{0}=0)J_{1}(\phi_{1}=0)\right)\otimes\\ &\quad\left(J_{2}(\phi_{4}=0)J_{2}\left(\phi_{5}=\frac{\pi}{4}\right)\right)CZ_{12}\\ &\quad\left(J_{1}\left(\phi_{2}=\frac{\pi}{4}\right)J_{1}(\phi_{3}=0)\right)\otimes\\ &\quad\left(J_{2}(\phi_{6}=0)J_{2}\left(\phi_{7}=-\frac{\pi}{4}\right)\right)CZ_{12}.\end{split}$ (5)
Using these gates, we can realize any quantum gates in VQAs. Additionally, the delegated computation adheres to the rules of MBQC, where measurement angles are adaptively chosen to move all Pauli byproducts to the leftmost position, resulting in $X^{x_{out}}Z^{z_{out}}U\ket{\Psi}_{R}$. By appropriately correcting the measurement results, the client ultimately obtains the desired state $U\ket{\psi}_{out}$. ∎
IV.3 Comparisons
Table 1: Comparison among different protocols for cloud-based VQAs.
Verifiable Client’s quantum capabilities Tolerance to channel loss Server’s resource consumption
Li et al.’s BFK-based protocol Li et al. (2021b) Yes Prepare qubits Yes $6w\cdot d$
Wang et al.’s BFK-based protocol Wang et al. (2022) Yes Measure qubits Yes $16w+$ 1
Shingu et al.’s protocol Shingu et al. (2022) No Measure qubits No $w+1$

### v3

@@ 3 The Proposed Protocol
We adopt the strategy outlined in [40], where the server performs measurements after coupling the ancillary qubits with the register qubits, effectively addressing the limitations of using Bell pairs to tolerate channel loss in Shingu et al.’s protocol. The client is required to perform measurements in the bases $\left\{M(-\frac{k\pi}{4})\mid k\in\{0,1,\ldots,7\}\right\}$. For verification, we use trap qubits to create dummy gates in the trap wires as the circuit is transformed into the universal gate patterns shown in Fig. 4(a). Additionally, we delegate encrypted measurements of the output register qubits to the server, making it difficult to identify the trap wires in the circuits.
@@
Figure 4: (a) The universal gate patterns: Composed of multiple blocks, each representing a gate pattern. (b) The gate pattern circuit: Consisting of 8 $J(\phi)$ operators and 2 $CZ$ gates. (c) Realization of the $J(\phi)$ operator: The server sends one half of a Bell pair to the client, who measures in the basis $M\left(-\frac{k\pi}{4}\right)$ with result $s_{0}$. The server then performs operations on the ancillary and register qubits, including a fixed coupling operation $(H_{R}\otimes I_{A})CX_{RA}$, where $CX$ denotes the controlled-X gate and $I$ is the identity gate, followed by measurements in the $Z$ basis or in the $M(\phi^{\prime})$ basis on the ancillary qubits with results $s_{1}$ to $s_{3}$.
The client delegates the $J(\phi)$ operators within the universal gate patterns shown in Fig. 4(a), while the server provides the $CZ$ gates. This process includes preparation, computation, verification, and parameter updating phases, as briefly illustrated in Fig. 5. The specific steps are as follows:
@@
Figure 5: Brief process of the proposed protocol: The server sends an ancillary qubit, as one half of a Bell pair, to the client, who measures it in the basis $M(-\frac{k\pi}{4})$. The client then sends the reception status back to the server, requesting a resend if the qubit is lost. The server performs encrypted measurements as instructed by the client and returns the results, enabling the client to verify the server’s honesty and calculate encrypted measurement angles and gradients for the optimizer.
@@ 1. B1-1:
  2. B1-2:
    The server prepares $N$ register qubits |$\ket{0}_{R}$, one ancillary qubit |$\ket{+}_{A}$, one ancillary qubit |$\ket{0}_{A}$ and two ancillary qubits $\frac{1}{\sqrt{2}}(\ket{00}_{A}+\ket{11}_{A})$ as a Bell pair. Meanwhile, the client chooses $32N/$ 3 register qubits as trap qubits for verification, which is optimal [6].
@@ 2. B1-2:
  1. B2-1:
    The server sends one qubit from the Bell pair to the client, who measures it in the basis $M\left(-\frac{k\pi}{4}\right)$, yielding the result $s_{0}$, where $k$ is randomly and uniformly selected from $\{0,1,\ldots,7\}$. The client generates a random bit recv_status to indicate the reception status, setting it to 0 if the qubit arrived successfully or to 1 if it was lost. This status is communicated to the server, who generates a new Bell pair and resends half of it to the client until the qubit arrives successfully. The other half of the Bell pair will eventually be in the state:
    $Z^{s_{0}}R_{Z}\left(\frac{k\pi}{4}\right)\ket{+}_{A}=\frac{1}{\sqrt{2}}\left[\ket{0}_{A}+\exp\left(i\left(\frac{k}{4}+s_{0}\right)\pi\right)\ket{1}_{A}\right].$ (5)
  2. B2-2:
    The server employs the transformed circuits $\{U^{(c)}_{\rm{AN}}\}_{c=1}^{G}$ using the $J(\phi)$ operator, as depicted in Fig. 4(c), to generate the trial wave functions $\left\{\left|\psi^{(c)}(\vec{\theta}[1])\right\rangle\right\}_{c=1}^{G}$. In each delegation of the $J(\phi)$ operator, the $Z$ measurement results $s_{1}$ and $s_{2}$ are first sent to the client. The client then uses $s_{1}$ to compute the encrypted measurement angle:
    $\phi^{\prime}=-\phi+(-1)^{s_{1}}\left(\frac{k}{4}+s_{0}\right)\pi+r\pi,$ (6)
    where $r\in\{0,1\}$ is a random bit chosen by the client. This encrypted angle $\phi^{\prime}$ is sent to the server, which measures in the basis $M(\phi^{\prime})$. The resulting measurement outcome $s_{3}$ is then sent to the client. The actual measurement angle $\phi$ is adaptively chosen based on prior Pauli byproducts to ensure determinism in MBQC.
  3. B2-3:
    In each transformed circuit, $32N/$ 3 of the wires are randomly designated as trap wires, where dummy operations are implemented. On half of these trap wires, the identity gate is applied $M$ times on each wire. On the remaining trap wires, a single Hadamard gate is randomly inserted on each wire, while the identity gate occupies the other $M-1$ positions.
  4. B2-4:
    The server measures the output register qubits, each with Pauli byproducts $\{X^{x_{j}}\}_{j=1}^{N}$ and $\{Z^{z_{j}}\}_{j=1}^{N}$, in the $Z$ basis. To measure in the $X$ basis, the client applies an additional Hadamard gate at the end of the original circuit before measuring in the $X$ basis. This transformation can be expressed as:
[…]
    where |$\ket{\Psi}_{R_{j}}$ is the $j$-th output register qubit of |$\ket{\Psi}_{R}$ for $j=1,\dots,N$. $\braket{X_{R_{j}}}$ is the expectation value of $X$ in the state $X^{z_{j}}Z^{x_{j}}\ket{\Psi}_{R_{j}}$, and $\braket{Z_{R_{j}}}$ is the expectation value of $Z$ in the state $X^{x_{j}}Z^{z_{j}}H\ket{\Psi}_{R_{j}}$.
  5. B2-5:
    The client compensates for the Pauli byproducts on all output register qubits by flipping the measurement results according to $\left\{x_{j}\right\}_{j=1}^{N}$ and $\left\{z_{j}\right\}_{j=1}^{N}$.
  6. B2-6:
@@
  1. B3-1:
    At the end of B2-3, the output register qubits consist of non-trap qubits and trap qubits in a random permutation designed by the client. The output state can be written as:
[…]
    where |$\ket{\psi}_{out}$ represents the non-trap qubits, and $\sigma_{q}$ is the Pauli byproduct operator. The subscript $R$ includes both non-trap qubits ($out$) and trap qubits ($T$), with $P$ as the permutation.
  2. B3-2:
    The client instructs the server to measure all desired trap qubits, |$\ket{0}_{T}^{\frac{N}{3}}$ in the $Z$ basis and |$\ket{+}_{T}^{\frac{N}{3}}$ in the $X$ basis, in B2-4. If any undesired output |$\ket{1}_{T}$ in the $Z$ basis or |$\ket{-}_{T}$ in the $X$ basis is obtained after compensating for the Pauli byproduct effect, the protocol is terminated.
@@ 2. B3-2:
  1. B4-1:
    The server and the client repeat B2 to B3 $\{\mathcal{R}^{(c)}\}_{c=1}^{G}$ times to derive the expectation values of $Z$ or $X$ in each output register qubit.
  2. B4-2:
    The client utilizes its optimizer to update the parameters according to Eq. 4, resulting in $\vec{\theta}[2]=(\theta_{1}[2],\cdots,\theta_{L}[2])^{T}$ for the next step.
  3. B4-3:
    The client calculates the measurement angles for circuits $\{U^{(c)}_{\rm{AN}}\}_{c=1}^{G}$ using $\vec{\theta}[2]$.
  4. B4-4:
    The client and server reiterate the above steps $(\mathcal{I}-2)$ times with $\{U^{(c)}_{\rm{AN}}\}_{c=1}^{G}$ and $\vec{\theta}[j]$. Based on the results from the $j$-th step, the client’s optimizer updates the parameters to $\vec{\theta}[j+1]$ for $j=2,3,\ldots,\mathcal{I}-1$.
4 Analysis
4.1 Verifiability
We use trap qubits to detect malicious operations on the output register qubits. The verifiability of the proposed protocol is demonstrated in Theorem I, which follows a method similar to that in [6], where further details can be found.
Theorem 1.
The probability of the client being tricked by the server is exponentially small.
@@ 4.1 Verifiability
If the server is malicious, it may deviate from the state:
$\rho=\sigma_{q}|\Psi\rangle_{R}\langle\Psi|_{R}\sigma^{\dagger}_{q}$ (9)
to an arbitrary state. However, due to the completely positive trace preserving (CPTP) map [11], this deviation can be detected as a random Pauli attack.
@@
We can calculate the probabilities that each operator in $\sigma_{\alpha}$ does not change any trap qubits. Suppose $\max(a,b,c)=a$. An $X$ operator that does not change any trap qubits will only act on |$\ket{+}_{T}$ and non-trap qubits, the number of which is $32N/$ 3. Thus, the probability is:
@@
The proposed protocol requires the client to send qubits, qubit reception statuses, and encrypted measurement angles, which seems to conflict with the no-signaling principle. However, the permutation $P$ is not transmitted, ensuring its secrecy. The server is only aware of measuring in the $Z$ basis, while the $X$ basis is implemented by delegating an additional Hadamard gate at the end of the original circuit. Consequently, the server cannot distinguish between trap and non-trap qubits, thereby further protecting $P$ without any leakage to the server.
After the client randomly selects a permutation $P$, the probability that $P^{\dagger}\sigma_{\alpha}P$ does not alter any trap qubits is at most $\left(\frac{2}{3}\right)^{\frac{|\alpha|}{3}}$. Consequently, the probability that the server deceives the client is at most:
$\prod_{c=1}^{G}\left(\frac{2}{3}\right)^{\frac{\mathcal{R}^{(c)}|\alpha|}{3}}=\left(\frac{2}{3}\right)^{\frac{\sum_{c=1}^{G}\mathcal{R}^{(c)}|\alpha|}{3}},$ (11)
where $G$ is the number of original and variant quantum circuits used to calculate gradients, and $\{\mathcal{R}^{(c)}\}_{c=1}^{G}$ is the set of circuit repetitions. This probability becomes exponentially small when $\sum_{c=1}^{G}\mathcal{R}^{(c)}$ is sufficiently large, ensuring the protocol’s verifiability.
4.2 Blindness and Correctness
@@
Theorem 2.
The proposed protocol guarantees input, output, and algorithm blindness.
@@
Input Blindness: After the client measures one qubit of the Bell pair, the remaining qubit is left in a maximally mixed state:
$\frac{1}{16}\sum_{s_{0}=0}^{1}\sum_{k=0}^{7}\left(Z^{s_{0}}R_{Z}\left(\frac{k\pi}{4}\right)\ket{+}_{A}\bra{+}_{A}R_{Z}^{\dagger}\left(\frac{k\pi}{4}\right)Z^{s_{0}\dagger}\right)=\frac{I}{2},$ (12)
with the value of $k$ hidden from the server.
Output Blindness: After applying the $J(\phi)$ operator, the register qubit is $X^{s_{3}\oplus r}HR_{Z}(\phi)\ket{\psi}_{R}$. Upon completion of the delegated circuit, the output register qubits are $X^{x_{out}}Z^{z_{out}}U\ket{\Psi}_{R}$, where $X^{x_{out}}Z^{z_{out}}$ are Pauli byproducts. Since the server cannot determine $r$, it cannot compensate for $X^{s_{3}\oplus r}$, and therefore cannot compensate for $X^{x_{out}}Z^{z_{out}}$, leaving the output qubits in a maximally mixed state.
Algorithm Blindness: When the client sends the encrypted measurement angle $\phi^{\prime}=-\phi+(-1)^{s_{1}}\left(\frac{k\pi}{4}+s_{0}\pi\right)+r\pi$ to the server, the true measurement angle $\phi$ remains concealed, as $k$ and $r$ are randomly chosen, and $s_{0}$ is never revealed.
@@
Theorem 3.
The proposed protocol ensures correctness throughout the computation.
@@
In each gate pattern, as depicted in Fig. 4(b), the server directly applies the $CZ$ gate, while the $J(\phi)$ operator is implemented using the circuit shown in Fig. 4(c). This procedure results in the state $HR_{Z}(\phi)\ket{\psi}_{R}$ with Pauli byproducts. Further details on this mathematical derivation are provided in  A.
The client subsequently adjusts the measurement results provided by the server to account for these Pauli byproducts, ensuring the correct realization of $J(\phi)$ operators within the universal gate patterns. Each gate pattern is parameterized by $\{\phi_{i}\}_{i=0}^{7}$, enabling the implementation of arbitrary single-qubit gates. For example, the operation $U_{1}\otimes U_{2}$ can be achieved by setting $\phi_{5}=\phi_{7}=0$ while appropriately selecting the remaining parameters. Here, $U_{1}$ and $U_{2}$ denote arbitrary single-qubit gates, which can be decomposed as $U_{1}=R_{Z}(\phi_{0})R_{x}(\phi_{1})R_{Z}(\phi_{2})$ and $U_{2}=R_{Z}(\phi_{4})R_{x}(\phi_{5})R_{Z}(\phi_{6})$, following Euler’s rotation theorem [37]. Additionally, the $CX$ gate can be realized within this pattern by setting $\phi_{2}=\phi_{5}=\frac{\pi}{2}$ and $\phi_{7}=-\frac{\pi}{2}$, while setting the others to 00.
Using these gates, any quantum gate in VQAs can be realized. Moreover, the delegated computation adheres to the rules of MBQC, where measurement angles are adaptively chosen to move all Pauli byproducts to the leftmost position, resulting in $X^{x_{\text{out}}}Z^{z_{\text{out}}}U\ket{\Psi}_{R}$. By appropriately correcting the measurement results, the client ultimately obtains the desired state $U\ket{\psi}_{\text{out}}$. ∎
Table 1: Comparison among different protocols for cloud-based VQAs
Verifiable Client’s quantum capabilities Tolerance to channel loss Server’s quantum resource consumption
Li et al.’s BFK-based protocol [27] Yes Prepare qubits Yes $6w\cdot d$
Wang et al.’s BFK-based protocol [28] Yes Measure qubits Yes $16w+$ 1
Shingu et al.’s protocol [30] No Measure qubits No $w+1$